# Adaptive Command: Real-Time Policy Adjustment via Language Models in StarCraft II

### v1

@@ 1. Introduction
Real-time strategy (RTS) games, exemplified by StarCraft II (SC2), present formidable challenges in AI research, demanding rapid tactical decisions and adaptive long-term planning. SC2’s complex gameplay, detailed in Section 9, encompasses resource management, base construction, and military command, serving as an ideal testbed for advanced AI systems. While previous efforts, such as DeepMind’s AlphaStar (Vinyals et al., 2019), have shown AI’s potential in mastering such games, recent advancements have begun to explore the integration of Large Language Models (LLMs) in this domain.
The rapid advancement of LLMs in reasoning, planning, and decision-making has opened new possibilities for their application in complex gaming environments. A notable contribution in this area is the TextStarCraft II project (Ma et al., 2024), which developed a specialized environment for assessing LLMs in real-time strategic scenarios within SC2. Their work demonstrated that fine-tuned LLMs could perform on par with Gold-level players in real-time matches, showcasing the potential of LLMs in strategic gameplay.
@@ 2. Related Work
StarCraft II Full Game AI: StarCraft AI research, initially focused on StarCraft I with developments like BiCNet (Peng et al., 2017) for multi-agent coordination, has significantly advanced in the StarCraft II era. The release of PySC2 (Vinyals et al., 2017) by DeepMind, coupled with Blizzard’s game replays, propelled this research field. A key breakthrough was AlphaStar (Vinyals et al., 2019), which achieved Grandmaster level and defeated top players, demonstrating the potential of RL in complex environments.
Subsequent research (Wang et al., 2021a; Christianos et al., 2021; Jiang et al., 2021; Hu et al., 2024) expanded upon these foundations. Mini-AlphaStar(Liu et al., 2021b) simplified input variables without compromising learning effectiveness. TG(Liu et al., 2021a) and HierNet-SC2(Liu et al., 2022) explored efficient RL strategies, with the latter bypassing supervised pre-training. AlphaStar Unplugged(Mathieu et al., 2021) represented a leap in offline RL using human replays. TStarBotsX(Han et al., 2020) and SCC(Wang et al., 2021b) furthered federated learning approaches, achieving notable success against master and grandmaster level players.
Recent advancements include DI-star [^1] https://github.com/opendilab/DI-star, which is accessible for home computer deployment, and ROA-Star(Huang et al., 2023), enhancing AlphaStar’s training framework with goal-conditioned exploiters and refined opponent modeling techniques. ROA-Star’s practical tests against professional players have shown impressive results, marking significant progress in real-time strategy AI.
Grounding Natural Language Commands to StarCraft II(Waytowich et al., 2019) made a pioneering contribution in this direction with their work on Grounding Natural Language Commands to StarCraft II Game States.
Building upon this integration of language and game mechanics, TextStarCraft II (Ma et al., 2024) further advanced the field by introducing a specialized environment for evaluating Large Language Models (LLMs) in real-time strategic scenarios within SC2. This work highlights the potential of language models in strategic gameplay and decision-making.
@@
LLM Agents and Benchmarks: The surge of large language models (LLMs) in late 2022, exemplified by models such as , GPT-3.5,GPT4 (OpenAI, 2023), and subsequent developments like LLaMA (Dubey et al., 2024) and Qwen (Bai et al., 2023; Yang et al., 2024), has significantly propelled research into LLM agents. Projects like React , Chain of Thought (Yao et al., 2022; Wei et al., 2022; Wang et al., 2022), and AutoGPT laid the groundwork for more sophisticated implementations. In gaming environments, initiatives such as GITM (Zhu et al., 2023) and Voyager (Wang et al., 2023) have demonstrated LLM agents’ adaptability to various tasks and open-world scenarios. Environments like TextWorld (Côté et al., 2019) and ALFWorld (Shridhar et al., 2020) have integrated text-based strategy and action execution. Recent advancements include Cradle (Tan et al., 2024), which introduces a General Computer Control setting for broader software interaction, and strategic discussion agents for games like One Night Ultimate Werewolf (Jin et al., 2024). Benchmarking platforms like AGENTBENCH (Liu et al., 2023) play a crucial role in evaluating these developments in comprehensive, open-ended contexts.
Human-AI Collaboration in Gaming: Recent advancements in Large Language Models (LLMs) have opened new avenues for human-AI collaboration in complex gaming environments(Feng et al., 2024). MindAgent (Gong et al., 2023) introduces a framework for evaluating AI planning and coordination in multi-agent gaming scenarios, while practical applications like the AI teammate in ”Naraka: Bladepoint” mobile game  [^2] https://fuxi.163.com/database/1179 demonstrate the potential of AI assistants in commercial games.
@@ 3.3. Game State Processing
We adopt the Chain of Summarization (CoS) method, originally introduced in the TextStarCraft II framework (Ma et al., 2024), for efficient game state processing. This method involves:
@@ 5. Experiment Design and Evaluation
[…]
A four-part figure showcasing the progression of a StarCraft II game using Adaptive Command. It illustrates the initial strategy, transition to air units, balanced army composition, and final victory. Below the images is a table showing examples of human input and corresponding LLM responses for each phase. Phase Human Input LLM Response Strategy Init. Phase (Fig. 2(a)) ”Given the opponent is using heavily armored units, what is the best counter-strategy?” Recommends Roach-Hydra composition for balance of survivability and damage against armored units. Adaptive Exec. Phase (Fig. 2(b), 2(c)) ”I want to play a sky army style” Suggests transitioning to Mutalisks and Corruptors to counter Protoss air units and control the map. ”I think we should also produce some ground army, right?” Advises maintaining a small ground force for versatility and defense against various Protoss ground units.

### v2

@@ 1. Introduction
Real-time strategy (RTS) games, exemplified by StarCraft II (SC2), present formidable challenges in AI research, demanding rapid tactical decisions and adaptive long-term planning. SC2’s complex gameplay, detailed in Section 9, encompasses resource management, base construction, and military command, serving as an ideal testbed for advanced AI systems. While previous efforts, such as DeepMind’s AlphaStar (alphastarnature), have shown AI’s potential in mastering such games, recent advancements have begun to explore the integration of Large Language Models (LLMs) in this domain.
The rapid advancement of LLMs in reasoning, planning, and decision-making has opened new possibilities for their application in complex gaming environments. A notable contribution in this area is the TextStarCraft II project (ma2024largelanguagemodelsplay), which developed a specialized environment for assessing LLMs in real-time strategic scenarios within SC2. Their work demonstrated that fine-tuned LLMs could perform on par with Gold-level players in real-time matches, showcasing the potential of LLMs in strategic gameplay.
@@ 2. Related Work
StarCraft II Full Game AI: StarCraft AI research, initially focused on StarCraft I with developments like BiCNet (peng2017multiagent) for multi-agent coordination, has significantly advanced in the StarCraft II era. The release of PySC2 (pysc2) by DeepMind, coupled with Blizzard’s game replays, propelled this research field. A key breakthrough was AlphaStar (alphastarnature), which achieved Grandmaster level and defeated top players, demonstrating the potential of RL in complex environments.
Subsequent research (wang2021scc; christianos2021scaling; jiang2021multi; hu2024learning) expanded upon these foundations. Mini-AlphaStar(liu2021introduction) simplified input variables without compromising learning effectiveness. TG(liu2021efficient) and HierNet-SC2(liu2022onefficient) explored efficient RL strategies, with the latter bypassing supervised pre-training. AlphaStar Unplugged(starcraft2unplugged) represented a leap in offline RL using human replays. TStarBotsX(Tstarbot-x) and SCC(scc) furthered federated learning approaches, achieving notable success against master and grandmaster level players.
Recent advancements include DI-star [^1] https://github.com/opendilab/DI-star, which is accessible for home computer deployment, and ROA-Star(ROA-Star), enhancing AlphaStar’s training framework with goal-conditioned exploiters and refined opponent modeling techniques. ROA-Star’s practical tests against professional players have shown impressive results, marking significant progress in real-time strategy AI.
Grounding Natural Language Commands to StarCraft II(groundingstar) made a pioneering contribution in this direction with their work on Grounding Natural Language Commands to StarCraft II Game States.
Building upon this integration of language and game mechanics, TextStarCraft II (ma2024largelanguagemodelsplay) further advanced the field by introducing a specialized environment for evaluating Large Language Models (LLMs) in real-time strategic scenarios within SC2. This work highlights the potential of language models in strategic gameplay and decision-making.
@@
LLM Agents and Benchmarks: The surge of large language models (LLMs) in late 2022, exemplified by models such as , GPT-3.5,GPT4 (gpt4_report), and subsequent developments like LLaMA (llama3) and Qwen (bai2023qwen; yang2024qwen2), has significantly propelled research into LLM agents. Projects like React , Chain of Thought (react; chainofthought; chainofthoughtsc), and AutoGPT laid the groundwork for more sophisticated implementations. In gaming environments, initiatives such as GITM (ghostintheminecraft) and Voyager (voyager) have demonstrated LLM agents’ adaptability to various tasks and open-world scenarios. Environments like TextWorld (cote2019textworld) and ALFWorld (shridhar2020alfworld) have integrated text-based strategy and action execution. Recent advancements include Cradle (cradle), which introduces a General Computer Control setting for broader software interaction, and strategic discussion agents for games like One Night Ultimate Werewolf (jinwerewolf). Benchmarking platforms like AGENTBENCH (agentbench) play a crucial role in evaluating these developments in comprehensive, open-ended contexts.
Human-AI Collaboration in Gaming: Recent advancements in Large Language Models (LLMs) have opened new avenues for human-AI collaboration in complex gaming environments(feng2024large). MindAgent (mindagnet) introduces a framework for evaluating AI planning and coordination in multi-agent gaming scenarios, while practical applications like the AI teammate in ”Naraka: Bladepoint” mobile game  [^2] https://fuxi.163.com/database/1179 demonstrate the potential of AI assistants in commercial games.
@@ 3.3. Game State Processing
We adopt the Chain of Summarization (CoS) method, originally introduced in the TextStarCraft II framework (ma2024largelanguagemodelsplay), for efficient game state processing. This method involves:
@@ 5. Experiment Design and Evaluation
A four-part figure showcasing the progression of a StarCraft II game using Adaptive Command. It illustrates the initial strategy, transition to air units, balanced army composition, and final victory. Below the images is a table showing examples of human input and corresponding LLM responses for each phase.
Phase Human Input LLM Response
Strategy Init. Phase (Fig. 2(a)) ”Given the opponent is using heavily armored units, what is the best counter-strategy?” Recommends Roach-Hydra composition for balance of survivability and damage against armored units.
Adaptive Exec. Phase (Fig. 2(b), 2(c)) ”I want to play a sky army style” Suggests transitioning to Mutalisks and Corruptors to counter Protoss air units and control the map.
”I think we should also produce some ground army, right?” Advises maintaining a small ground force for versatility and defense against various Protoss ground units.
@@ 9.2. Competitive Scene and Research Platform
10. Acknowledgements
This work was supported by the National Science and Technology Major Project 2022ZD0116404.